OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3). y = x − 1 x= 3over2 y = x + 3 x= 3over2 y = −x − 1 -x=-2over3 y = −x + 3 -x= -2over3

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

okay. the slope of the line is the opposite reciprocal of the line given so it would become -1

OpenStudy (anonymous):

then all you would have to do is take the point -6,3 and plug it into the equation y=-1x+b

OpenStudy (anonymous):

does it make sense so far?

OpenStudy (anonymous):

Perpendecular <means algebraically> Slope2 * Slope1 = -1

OpenStudy (anonymous):

so you would do -6=-3+b b=-3 the equation would be y=-x-3

OpenStudy (anonymous):

thank you!